OpenStudy (anonymous):

Hello, i need help with my Algebra. Here's the homework that my teacher gave me. Write the following expression/equations in words: a) X - 7=(y+3) b) P=2w+2l c) 12+3(x-4) d)xy-4+12

OpenStudy (anonymous):

Hello, i need help with my Algebra. Here's the homework that my teacher gave me. Write the following expression/equations in words: a) X - 7=(y+3) b) P=2w+2l c) 12+3(x-4) d)xy-4+12

OpenStudy (anonymous):

X minus seven equals y plus three P equals two times w plus twenty one Twelve plus three times x minus three times four X times y minus four plus twelve

OpenStudy (anonymous):

Aww thanks brandon376 i apreciate your help. :)